Part a The Deli University Gift Shop purchases sweatshirts emblazoned with the school name and logo...

Part a

The Deli University Gift Shop purchases sweatshirts emblazoned with the school name and logo from a vendor in Spain Town at a cost of $2,000 each. The annual holding cost for a sweatshirt is calculated as 1.5% of the purchase cost. It costs the Gift Shop $500 to place a single order. The Gift Shop manager estimates that 900 sweatshirts will be sold during each month of the upcoming academic year.

i) Determine the highest number of shirts that should be purchased by the Gift Shop in order to minimize stock administration costs.

ii) What is the number of orders to be placed each year?
Answer #1

Answer i

EOQ = Sqrt(2*Annual usage*ordering cost per order/carrying cost per unit for a year)

= Sqrt(2*900*12*500/(1.5%*2,000))

= Sqrt(10,800,000/30)

= Sqrt(360,000)

= 600 shirts per order

Answer ii

No. of orders to placed in a year = shirts sold in a year/EOQ

= 900*12/600

= 18 orders

In case of any doubt, please comment.